40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3917     Entry time: Sun Nov 14 16:40:46 2010
Author: Jenne 
Type: Update 
Category: Treasure 
Subject: Activities related to OSEM measurement 

[Valera, Jenne]

We pondered the idea of clamping the PRM optic to measure the OSEM noise.  So we opened up the BS tank to give this a try.  We rediscovered that Jenne is too short to reach the other side of the PRM tower, so we couldn't fully clamp the optic (when is Jaime coming again? He's kind of tall...)  If we only did the back 2 EQ stops, the optic would still be able to rock, and thus defeat the purpose of clamping anyway.  So we didn't go for it. 

While we were in there we saw that the SRM OSEMs were just hanging out on the table, and decided to go with them.  See Valera's elog for details on our measurement.  We closed up the tank without making any changes to anything.

In other news, we still need to figure out how to change up the connectors to get those OSEMs over to the ITM table.  This needs to happen pretty soonish. 

ELOG V3.1.3-